OpenStudy (anonymous):

rename 28 percent as a decimal and fraction in its simplest form

OpenStudy (aravindg):

28%=28/100=14/50=7/25 simplest fraction^^ 28/100=.28 in decimal

OpenStudy (aravindg):

please click best response

OpenStudy (anonymous):

For sure!! Thanks again dude